# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 21 Apr 2026 v4

@@ 6 Hausdorff Dimension, Moments, and Fourier Structure of $G$
$\mathbb{E}[X^{k}]\;=\;2^{-k}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!k}\right).$ (5)
@@
Since $Y=\log 2+\sum_{\ell}Z_{\ell}$ with the $Z_{\ell}\geq 0$ independent (Lemma 10), the partial sums $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}\nearrow S:=\sum_{\ell}Z_{\ell}$ give $e^{-kS_{L}}\searrow e^{-kS}$. Since $0\leq e^{-kS_{L}}\leq 1$ for all $L$, by dominated convergence,
@@
Since $\mathbb{E}[X^{k}]=2^{-k}\,\mathbb{E}[e^{-kS}]$, this gives (5).
@@
Corollary 5 (First and second moments).
  1. (i)
    $\mathbb{E}[X]=\frac{1}{2}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{2}-\ell-1}{\ell(\ell-1)}.$
  2. (ii)
    $\mathbb{E}[X^{2}]=\frac{1}{4}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{3}-\ell^{2}-2\ell+1}{\ell^{2}(\ell-1)}.$
  3. (iii)
    $\operatorname{Var}(X)=\mathbb{E}[X^{2}]-(\mathbb{E}[X])^{2}$.
@@
Immediate from (5) with $k=1$ and $k=2$. ∎
Theorem 7 (Moment determinacy).
The moment sequence $\{\mathbb{E}[X^{k}]\}_{k=1}^{\infty}$ uniquely determines $G$.
Proof.
Since $\operatorname{supp}(G)=[0,\tfrac{1}{2}]$ (Theorem 4), the random variable $X$ is supported on the compact interval $0,\tfrac{1}{2}]$. By the Weierstrass approximation theorem, polynomials are uniformly dense in $C([0,\tfrac{1}{2}])$, so the moment sequence $\{\mathbb{E}[X^{k}]\}_{k=1}^{\infty}$ determines the distribution of $X$ uniquely. ∎
Theorem 8 (Analytic continuation).
The function
$M(s)\;:=\;\mathbb{E}[X^{s}]\;=\;2^{-s}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}F_{s}(\ell),\qquad F_{s}(\ell):=\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!s},$
extends to an analytic non-vanishing function on $\{s\in\mathbb{C}:\operatorname{Re}(s)>0\}$, with the Euler product converging uniformly on compact subsets.
Proof.
@@
We first establish the product formula for complex $s$ with $\sigma:=\operatorname{Re}(s)>0$. By Lemma 10, $Y=\log 2+\sum_{\ell\geq 3}Z_{\ell}$ with independent $Z_{\ell}\geq 0$, and the series converges almost surely since $\sum_{\ell}\mathbb{E}[Z_{\ell}]<\infty$. Set $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}$. Since $Z_{\ell}\geq 0$ and $\sigma>0$,
$|e^{-sS_{L}}|=e^{-\sigma S_{L}}\leq 1$
@@
for every $L$. As $S_{L}\to S:=\sum_{\ell}Z_{\ell}$ a.s., dominated convergence gives $\mathbb{E}[e^{-sS_{L}}]\to\mathbb{E}[e^{-sS}]$. Independence factors the finite partial products: $\mathbb{E}[e^{-sS_{L}}]=\prod_{3\leq\ell\leq L}\mathbb{E}[e^{-sZ_{\ell}}]=\prod_{3\leq\ell\leq L}F_{s}(\ell)$. Therefore $\mathbb{E}[X^{s}]=2^{-s}\,\mathbb{E}[e^{-sS}]=2^{-s}\prod_{\ell\geq 3}F_{s}(\ell)$, extending the real-variable identity of Theorem 6 to the half-plane $\operatorname{Re}(s)>0$.
@@
It remains to verify analyticity and non-vanishing. Write $w_{\ell}:=\log(\ell/(\ell-1))$ and $F_{s}(\ell)=1+a_{\ell}(s)$ with $a_{\ell}(s)=(e^{-sw_{\ell}}-1)/(\ell-1)$. The bound $|e^{z}-1|\leq|z|e^{|z|}$ gives $|a_{\ell}(s)|\leq|s|\,w_{\ell}\,e^{|s|w_{\ell}}/(\ell-1)$. Since $w_{\ell}=1/\ell+O(1/\ell^{2})$, on any compact $K\subset\{\operatorname{Re}(s)>0\}$ with $|s|\leq R$ we get $|a_{\ell}(s)|\leq C_{K}/\ell^{2}$ for all large $\ell$. Since $\sum\ell^{-2}<\infty$, Weierstrass’s theorem on products of analytic functions (see e.g. [8, Ch. VII]) gives that the product converges uniformly on compact subsets to an analytic function. As $2^{-s}$ is entire, $M(s)$ is analytic on $\{\operatorname{Re}(s)>0\}$.
@@
For non-vanishing with $\sigma:=\operatorname{Re}(s)>0$: each factor $F_{s}(\ell)\neq 0$ since $|F_{s}(\ell)|\geq\frac{\ell-2}{\ell-1}-\frac{e^{-\sigma w_{\ell}}}{\ell-1}$ and $e^{-\sigma w_{\ell}}<1$: at $\ell=3$, $|F_{s}(3)|\geq(1-(2/3)^{\sigma})/2>0$; at $\ell\geq 5$, $|F_{s}(\ell)|\geq(\ell-3)/(\ell-1)>0$. To conclude that the infinite product is nonzero, fix a compact $K\subset\{\operatorname{Re}(s)>0\}$. Choose $L$ so that $|a_{\ell}(s)|<1/2$ for all $\ell\geq L$ and $s\in K$ (possible since $|a_{\ell}(s)|\leq C_{K}/\ell^{2}$). For $\ell\geq L$ the principal branch $\log(1+a_{\ell}(s))$ is analytic, and $|\!\log(1+a_{\ell}(s))|\leq 2|a_{\ell}(s)|\leq 2C_{K}/\ell^{2}$, so $\sum_{\ell\geq L}\log(1+a_{\ell}(s))$ converges uniformly on $K$ and the tail product equals $\exp\!\bigl(\sum_{\ell\geq L}\log(1+a_{\ell}(s))\bigr)\neq 0$. The finitely many factors with $\ell<L$ are each nonzero, so $M(s)\neq 0$. ∎
Proposition 5 (Squared modulus of the characteristic function).
@@
By Lemma 10, $|\widehat{\mu_{f}}(\tau)|^{2}=\prod_{\ell\geq 3}|\widehat{\nu_{\ell}}(\tau)|^{2}$. With $\alpha=\tau\log(\ell/(\ell-1))$ and $\widehat{\nu_{\ell}}(\tau)=\frac{\ell-2}{\ell-1}+\frac{e^{i\alpha}}{\ell-1}$,
@@
using $(\ell-2)^{2}+2(\ell-2)+1=(\ell-1)^{2}$. ∎
Proposition 6 (Fourier non-decay implies singularity).
If $\limsup_{|\tau|\to\infty}|\widehat{\mu_{f}}(\tau)|>0$, then $\mu_{f}$ is purely singular, and consequently $G$ is purely singular.
Proof.
By Lemma 6, $\mu_{f}$ is continuous and of pure type: either absolutely continuous or purely singular. The Riemann–Lebesgue lemma gives: if $\mu_{f}$ is absolutely continuous ($\mu_{f}=h\,\mathrm{d}t$, $h\in L^{1}$), then $\widehat{\mu_{f}}(\tau)=\widehat{h}(\tau)\to 0$. Contrapositively, $\limsup|\widehat{\mu_{f}}(\tau)|>0$ implies $\mu_{f}$ is not absolutely continuous, hence purely singular. For the conclusion about $G$: by Corollary 2, $X\overset{d}{=}e^{-Y}$ where $Y\sim\mu_{f}$, and $G$ is the distribution function of $X$. The map $T(u)=e^{-u}$ is locally Lipschitz on $0,\infty)$, so the push-forward $T_{\#}\mu_{f}$ preserves Lebesgue-null sets. Hence pure singularity of $\mu_{f}$ implies pure singularity of $G$ (cf. the proof of Corollary 2). ∎
Figure 3: Squared modulus $|\widehat{\mu_{f}}(\tau)|^{2}$ computed via the Euler product of Proposition 5 using all primes $\ell\leq 10^{7}$. The top panel uses a linear scale and the bottom panel a log-log scale. The function oscillates but its local averages decrease across successive decades of $\tau$.
Remark 2 (Numerical evidence on Fourier decay).
We evaluated $|\widehat{\mu_{f}}(\tau)|^{2}$ via the Euler product of Proposition 5 with all primes $\ell\leq 10^{7}$. Figure 3 displays the result for $\tau\in[1,10^{6}]$. The function oscillates wildly, with local maxima near $0.14$ persisting at moderately large $\tau$. However, the mean value within each decade decreases steadily: approximately $0.06$ for $\tau\in[10,10^{2}]$, $0.03$ for $10^{2},10^{3}]$, $0.02$ for $10^{3},10^{4}]$, $0.01$ for $10^{4},10^{5}]$, and $0.008$ for $10^{5},10^{6}]$. A power-law fit to the envelope of local maxima yields $|\widehat{\mu_{f}}(\tau)|^{2}\approx\tau^{-0.19}$. This is consistent with decay to 0, which would mean $\mu_{f}$ is a Rajchman measure; establishing this rigorously remains open (Section 8).
6.1 Endpoint asymptotics at $1/2$
The value $1/2$ is the supremum of the support of $G$ but is never attained (Remark 1). The rate at which $G$ approaches its endpoint governs, in particular, the density of primes $p$ for which the rejection-sampling overhead $1/c(p)$ is close to its minimum value of $2$. We now determine this rate precisely.
Proposition 7 (Endpoint asymptotic at $1/2$).
Let $G$ be the limiting distribution function from Theorem 4, and let $X$ be a random variable with distribution $G$. Define
@@ 6.1 Endpoint asymptotics at $1/2$
Set
$M(y):=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\frac{\ell-2}{\ell-1}.$
For $u>0$ one has
@@
$\mathbb{P}(S\leq t)\leq M\!\bigl((2t)^{-1}\bigr).$ (6)
@@
$\mathbb{P}(S\leq t)\geq M(z_{t})\,\mathbb{P}\!\left(\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}\leq t\right).$ (7)
@@
$M(y)=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{\ell}\right)\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{(\ell-1)^{2}}\right).$
The second product converges to $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$, while Theorem 1 gives
@@
$M(y)\sim\frac{\kappa}{\log y}\qquad(y\to\infty).$
Since $\log((2t)^{-1})\sim\log z_{t}\sim\log(1/t)$ as $t\downarrow 0$, (6) and (8) imply
@@
Since $\mu_{f}$ is continuous by Proposition 3, the law of $S=Y-\log 2$ is also continuous, so $\mathbb{P}(S<t)=\mathbb{P}(S\leq t)$. Therefore
@@
Remark 3 (Twin prime constant).
The constant $\kappa$ equals $C_{2}\,e^{-\gamma}$, where $C_{2}=2\prod_{p\geq 3}(1-(p-1)^{-2})\approx 1.3203$ is the twin prime constant of Hardy and Littlewood [22] (denote $2C_{2}$ therein). This follows from the factorisation $(\ell-2)/(\ell-1)=(1-1/\ell)(1-(\ell-1)^{-2})$: the first factor produces the Mertens asymptotic $2e^{-\gamma}/\!\log y$, while the product of the second converges to $C_{2}/2$.
Corollary 6 (Endpoint modulus at $1/2$).
@@
The two-sided order follows immediately from Proposition 7. If for some $\alpha>0$ one had
@@
then Proposition 7 would force
@@
7 Explicit Lower Bounds via the Number of Distinct Prime Factors
Theorem 9 (Explicit lower bound on $c(p)$ via the number of distinct prime factors).

### 06 May 2026 v5

@@ 6.2. Moments and Mellin transform
(6.1) $\mathbb{E}[X^{k}]\;=\;2^{-k}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!k}\right).$
@@
Since $Y=\log 2+\sum_{\ell}Z_{\ell}$ with the $Z_{\ell}\geq 0$ independent (Lemma 6.1), the partial sums $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}\nearrow S:=\sum_{\ell}Z_{\ell}$ give $e^{-kS_{L}}\searrow e^{-kS}$. Since $0\leq e^{-kS_{L}}\leq 1$ for all $L$, by dominated convergence,
@@
Since $\mathbb{E}[X^{k}]=2^{-k}\,\mathbb{E}[e^{-kS}]$, this gives (6.1).
@@
For example, substituting $k=1$ and $k=2$ into (6.1) gives
$\mathbb{E}[X]=\frac{1}{2}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{2}-\ell-1}{\ell(\ell-1)},\qquad\mathbb{E}[X^{2}]=\frac{1}{4}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{3}-\ell^{2}-2\ell+1}{\ell^{2}(\ell-1)}.$
Since $X$ is supported on the compact interval $0,\tfrac{1}{2}]$ (Theorem 5.5), the Weierstrass approximation theorem implies that the moment sequence $\{\mathbb{E}[X^{k}]\}_{k=1}^{\infty}$ determines $G$ uniquely.
Theorem 6.4 (Analytic continuation).
The function
$M(s)\;:=\;2^{-s}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}F_{s}(\ell),\qquad F_{s}(\ell):=\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!s},$
is an entire function of $s$, with the Euler product converging uniformly on compact subsets of $\mathbb{C}$. Moreover, $M(s)=\mathbb{E}[X^{s}]$ for every $s\in\mathbb{C}$, and $M(s)\neq 0$ whenever $\operatorname{Re}(s)>0$.
@@
We first establish the product formula $\mathbb{E}[X^{s}]=M(s)$ for every $s\in\mathbb{C}$. Write $\sigma:=\operatorname{Re}(s)$. By Lemma 6.1, $Y=\log 2+\sum_{\ell\geq 3}Z_{\ell}$ with independent $Z_{\ell}\geq 0$, and the series converges almost surely since $\sum_{\ell}\mathbb{E}[Z_{\ell}]<\infty$. Set $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}$ and $S:=\sum_{\ell}Z_{\ell}$. Since $Z_{\ell}\geq 0$, the partial sums are monotone: $S_{L}\leq S$ for every $L$. Hence
$|e^{-sS_{L}}|=e^{-\sigma S_{L}}\leq e^{|\sigma|S}$
@@
for every $L$. The dominator $e^{|\sigma|S}$ is integrable: independence gives $\mathbb{E}[e^{|\sigma|S}]=\prod_{\ell\geq 3}\mathbb{E}[e^{|\sigma|Z_{\ell}}]$, and each factor equals $\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\bigl(\frac{\ell}{\ell-1}\bigr)^{|\sigma|}=1+O_{|\sigma|}(1/\ell^{2})$, so the product converges. As $S_{L}\to S$ a.s., dominated convergence gives $\mathbb{E}[e^{-sS_{L}}]\to\mathbb{E}[e^{-sS}]$. Independence factors the finite partial products: $\mathbb{E}[e^{-sS_{L}}]=\prod_{3\leq\ell\leq L}\mathbb{E}[e^{-sZ_{\ell}}]=\prod_{3\leq\ell\leq L}F_{s}(\ell)$. Therefore $\mathbb{E}[X^{s}]=2^{-s}\,\mathbb{E}[e^{-sS}]=2^{-s}\prod_{\ell\geq 3}F_{s}(\ell)=M(s)$ for every $s\in\mathbb{C}$, extending the real-variable identity of Theorem 6.3. In particular, all negative moments $\mathbb{E}[X^{-\sigma}]$ ($\sigma>0$) are finite, a non-trivial consequence of the integrability of the dominator established above.
@@
It remains to verify that $M(s)$ is entire and non-vanishing on $\{\operatorname{Re}(s)>0\}$. Write $w_{\ell}:=\log(\ell/(\ell-1))$ and $F_{s}(\ell)=1+a_{\ell}(s)$ with $a_{\ell}(s)=(e^{-sw_{\ell}}-1)/(\ell-1)$. The bound $|e^{z}-1|\leq|z|e^{|z|}$ gives $|a_{\ell}(s)|\leq|s|\,w_{\ell}\,e^{|s|w_{\ell}}/(\ell-1)$. Since $w_{\ell}=1/\ell+O(1/\ell^{2})$, on any compact $K\subset\mathbb{C}$ with $|s|\leq R$ we get $|a_{\ell}(s)|\leq C_{R}/\ell^{2}$ for all large $\ell$. Since $\sum\ell^{-2}<\infty$, Weierstrass’s theorem on products of analytic functions (see e.g. [5, Ch. VII]) gives that the product converges uniformly on compact subsets of $\mathbb{C}$ to an analytic function. Each $F_{s}(\ell)$ is entire in $s$, and $2^{-s}$ is entire, so $M(s)$ is entire.
@@
For non-vanishing with $\sigma:=\operatorname{Re}(s)>0$: each factor $F_{s}(\ell)\neq 0$ since $|F_{s}(\ell)|\geq\frac{\ell-2}{\ell-1}-\frac{e^{-\sigma w_{\ell}}}{\ell-1}$ and $e^{-\sigma w_{\ell}}<1$: at $\ell=3$, $|F_{s}(3)|\geq(1-(2/3)^{\sigma})/2>0$; at $\ell\geq 5$, $|F_{s}(\ell)|\geq(\ell-3)/(\ell-1)>0$. To conclude that the infinite product is nonzero, fix a compact $K\subset\{\operatorname{Re}(s)>0\}$. Choose $L$ so that $|a_{\ell}(s)|<1/2$ for all $\ell\geq L$ and $s\in K$ (possible since $|a_{\ell}(s)|\leq C_{R}/\ell^{2}$). For $\ell\geq L$ the principal branch $\log(1+a_{\ell}(s))$ is analytic, and $|\!\log(1+a_{\ell}(s))|\leq 2|a_{\ell}(s)|\leq 2C_{R}/\ell^{2}$, so $\sum_{\ell\geq L}\log(1+a_{\ell}(s))$ converges uniformly on $K$ and the tail product equals $\exp\!\bigl(\sum_{\ell\geq L}\log(1+a_{\ell}(s))\bigr)\neq 0$. The finitely many factors with $\ell<L$ are each nonzero, so $M(s)\neq 0$. ∎
Remark 6.5 (Zeros on the imaginary axis).
The non-vanishing assertion in Theorem 6.4 is sharp: $M(s)$ has zeros on the imaginary axis. Indeed, $F_{s}(3)=\tfrac{1}{2}+\tfrac{1}{2}(2/3)^{s}$ vanishes at $s=i\pi(2k+1)/\!\log(3/2)$ for every $k\in\mathbb{Z}$. For $\ell\geq 5$, one checks $|F_{it}(\ell)|\geq(\ell-3)/(\ell-1)>0$ for all $t\in\mathbb{R}$, so every zero of $M$ on the imaginary axis arises from the factor at $\ell=3$.
6.3. Fourier structure
Proposition 6.6 (Squared modulus of the characteristic function).
@@ 6.3. Fourier structure
By Lemma 6.1, $|\widehat{\mu_{f}}(\tau)|^{2}=\prod_{\ell\geq 3}|\widehat{\nu_{\ell}}(\tau)|^{2}$. With $\alpha=\tau\log(\ell/(\ell-1))$ and $\widehat{\nu_{\ell}}(\tau)=\frac{\ell-2}{\ell-1}+\frac{e^{i\alpha}}{\ell-1}$,
@@
using $(\ell-2)^{2}+2(\ell-2)+1=(\ell-1)^{2}$. Each factor lies in $[0,1]$, so the infinite product converges in $[0,1]$; the question of whether it tends to zero as $|\tau|\to\infty$ is addressed in Section 9. ∎
Remark 6.7 (Fourier non-decay and singularity).
If $\limsup_{|\tau|\to\infty}|\widehat{\mu_{f}}(\tau)|>0$, then $\mu_{f}$ is purely singular by the Riemann–Lebesgue lemma and the pure-type dichotomy (Theorem 5.9); the Lipschitz push-forward under $u\mapsto e^{-u}$ then gives singularity of $G$. Thus Fourier non-decay would provide an independent, Fourier-analytic proof of singularity.
Remark 6.8 (Numerical evidence on Fourier decay).
Numerical evaluation of $|\widehat{\mu_{f}}(\tau)|^{2}$ via the Euler product of Proposition 6.6, with the product truncated at all primes $\ell\leq 10^{7}$ (664,579664{,}579 factors), shows that the function oscillates, with local maxima near $0.14$ persisting at moderately large $\tau$. Convergence was verified by comparing partial products up to $10^{6}$, $5\times 10^{6}$, and $10^{7}$: the relative change from the second to the third cutoff is below $10^{-4}$ throughout the range tested. The arithmetic mean within each decade (over $10^{3}$ logarithmically spaced points in $1,10^{6}]$) decreases steadily: approximately $0.06$ for $\tau\in[10,10^{2}]$, declining to $0.008$ for $10^{5},10^{6}]$. This monotone decline is consistent with $|\widehat{\mu_{f}}(\tau)|^{2}\to 0$, which would mean $\mu_{f}$ is a Rajchman measure—a continuous measure whose Fourier transform vanishes at infinity (see Lyons [22]); establishing this rigorously remains open (Section 9).
Figure 3. Squared modulus $|\widehat{\mu_{f}}(\tau)|^{2}$ computed via the Euler product of Proposition 6.6 using all primes $\ell\leq 10^{7}$. The top panel uses a linear scale and the bottom panel a log-log scale. The function oscillates but its local averages decrease across successive decades of $\tau$.
6.4. Endpoint asymptotics at $1/2$
The value $1/2$ is the right endpoint of the support of $G$ and carries zero mass; in particular, $c(p)<1/2$ for every prime $p$ (Remark 5.16), even though $\varphi(p-1)/(p-1)=1/2$ for Fermat primes. The rate at which $G$ approaches its endpoint governs, in particular, the density of primes $p$ for which the reciprocal $1/c(p)$ is close to its limiting infimum $2$. We now determine this rate precisely.
Proposition 6.9 (Endpoint asymptotic at $1/2$).
Let $G$ be the limiting distribution function from Theorem 5.5, and let $X$ be a random variable with distribution $G$. Define
@@ 6.4. Endpoint asymptotics at $1/2$
Recall from Lemma 5.10 the notation $Q(y):=\prod_{3\leq\ell\leq y,\,\ell\textup{ prime}}(\ell-2)/(\ell-1)$. For $u>0$ one has
@@
(6.2) $\mathbb{P}(S\leq t)\leq Q\!\bigl((2t)^{-1}\bigr).$
@@
(6.3) $\mathbb{P}(S\leq t)\geq Q(z_{t})\,\mathbb{P}\!\left(\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}\leq t\right).$
@@
$Q(y)=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{\ell}\right)\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{(\ell-1)^{2}}\right).$
The second product converges to $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$, while Theorem 3.1 gives
@@
$Q(y)\sim\frac{\kappa}{\log y}\qquad(y\to\infty).$
Since $\log((2t)^{-1})\sim\log z_{t}\sim\log(1/t)$ as $t\downarrow 0$, (6.2) and (6.4) imply
@@
Since $\mu_{f}$ is continuous by Proposition 5.13, the law of $S=Y-\log 2$ is also continuous, so $\mathbb{P}(S<t)=\mathbb{P}(S\leq t)$. Therefore
@@
Corollary 6.10 (Twin-prime constant).
The constant $\kappa$ equals $\mathfrak{S}_{2}\,e^{-\gamma}$, where
$\mathfrak{S}_{2}\;:=\;2\prod_{p\geq 3}\Bigl(1-\frac{1}{(p-1)^{2}}\Bigr)\;\approx\;1.3203$
is the singular series for twin primes in the sense of Hardy and Littlewood [15]. Hardy and Littlewood write $2C_{2}$ for this quantity; we use $\mathfrak{S}_{2}$ to avoid collision with the convention that reserves $C_{2}\approx 0.6602$ for $\prod_{p\geq 3}p(p-2)/(p-1)^{2}$.
Proof.
The identity $\kappa=\mathfrak{S}_{2}\,e^{-\gamma}$ follows from the factorisation $(\ell-2)/(\ell-1)=(1-1/\ell)(1-(\ell-1)^{-2})$: the first factor produces the Mertens asymptotic $2e^{-\gamma}/\!\log y$, while the product of the second converges to $\mathfrak{S}_{2}/2$. ∎
Corollary 6.11 (Endpoint modulus at $1/2$).
@@
The two-sided order follows immediately from Proposition 6.9. If for some $\alpha>0$ one had
@@
then Proposition 6.9 would force
@@
7. Explicit Lower Bounds via the Number of Distinct Prime Factors
The explicit bounds below are deterministic and hold for every prime $p\geq 3$; their strength depends on $\omega(p-1)$. By the shifted-prime analogue of the Erdős–Kac law [12] (proved in [16, Theorem 6]), $\omega(p-1)$ is typically of order $\log\log p$. Since the $K$-th prime satisfies $p_{K}\sim K\log K$ by the prime number theorem, Mertens’ third theorem gives $B(\omega(p-1))\asymp 1/\!\log\log\log p$ for a typical prime. The sharper scale $1/\!\log\log p$ arises in Corollary 7.2 and Corollary 7.3 from the primorial comparison $\omega(p-1)\leq m(p-1)$ together with $p_{m(p-1)}\sim\log p$. Actual divisibility of p1p-1 by all primes up to $p_{m(p-1)}$ occurs only in the Linnik/primorial-progression construction used for the matching upper bound and sharpness arguments (Proposition 4.2, Theorem 7.4). For primes with unusually many prime factors in p1p-1, the bound is correspondingly weaker.
Theorem 7.1 (Explicit lower bound on $c(p)$ via the number of distinct prime factors).